I have the answer but I am unsure how they got the answer.
The question is "Convert 85% to a fraction" the answer I have is 17/20.
how do I figure this out?

Answer by Mathtut(3670)

.85=85/100=17/20 .....so you are correct
:
remember % is just a part of 100
:
so 33% is just 33/100
:
and 133% is 133/100